The histogram below shows the same data that are shown in one of the four boxplots. Given that the data has mean -2s, which boxplot goes with the histogram? Please justify your answer. 10 15253 35 . A population is normally distributed with a mean of 55 and a standard deviation of 4. Using the Empirical Rule, approximately what percentage of data values fall between 51 and 59? Please show your calculations. A 68% 1.15% с 95% D 99.7% You are given the formula P120<x<80)-95% . Does this formula represent (please justify your answer) the middle 95% of the standard normal distribution? the middle 95% of a normal distribution? c. the 95h percentile of a standard normal distribution? o. the 95th percentile of the normal distribution? TRUE or FALSE: The purpose of the regression line is to determine the strength of the correlation between two quantitative variables. Please justify your answer . Which of the following is NOT a possible value of the correlation coefficient? +1.01 - 1.00 +1.00 B. C D. +0.05
Answer #1

H)

MeanM=55

Standard deviation S =4

X1=51

X2=59

Z1=(X1-M)/S

Z1= 51-55/4

Z1=-1

Z2 =(X2-M)/S

Z2=59-55/4

Z2=+1

From standard normal distribution table

Probability=0.3413+0.3413

=0.6826 or 68.26%

K)

Correlation coefficient lies between -1 to +1

Here, +1.01 out of boundary

So,it is not a possible value.